### v5

@@ III Simplex path integral
where $\operatorname{tr}\left(\cdot\right)$ denotes the trace and parameter $\tau$ serves as the inverse of a finite temperature. Please see Fig. 2(b) for instances. The Boltzmann constant is assumed as unitary for simplicity. The partition function, $\sum_{\lambda^{\left(n\right)}}\exp\left(-\tau\lambda^{\left(n\right)}\right)$, is proportional to the average return probability of random walks within a time scale $\tau$ [42]. See Figs. 2(b-c) for illustrations.
@@
where notion $\hbar$ denotes the reduced Planck constant (i.e., or named as the Dirac constant) [3, 4]. Eq. (14) is closely related to the imaginary time evolution governed by operator $\mathbf{L}^{\left(n\right)}$ in the infinite-dimensional path space. Specifically, the probability amplitude for the system to evolve from $\big{|}\mathbf{x}_{0}\big{\rangle}$ to $\big{|}\mathbf{x}\big{\rangle}$ during a period of $\tau$ conveyed by the propagator in Eq. (13) can be measured by summing over all possible evolution paths connecting between these two states (i.e., the discrete counterpart of path integrals) in Eq. (14). This is the reason why the $n$-order network topology conveyed by $\mathbf{L}^{\left(n\right)}$ intrinsically shapes path integrals. See Fig. 2(c) for illustrations.
@@ IV Simplex renormalization group
The connection between Eqs. (8-16) and the moment space naturally inspires us to consider the generalization of renormalization groups. The freedom degrees corresponding to short-range interactions (e.g., fast diffusion within the clusters of strongly correlated units) can be safely coarse grained without significant information loss, which is consistent with our interpretation of Eq. (10). Here we develop a framework to generalize renormalization groups to high-order interactions, where we also propose possible ways to overcome the limitations of the Laplacian renormalization group (LRG) [10].
@@ IV.1 Renormalization procedure
Figure 3: Illustrations of the simplex renormalization group (SRG). (a) The conceptual illustration of the SRG is presented. For the sake of legibility, the simplicial complexes of different orders are not elaborated and are abstractly represented by dashed lines with different colors. Given a system where multi-order interactions coexist, the SRG can be applied to renormalize the system on the $p$-order based on the structure and dynamics associated with $q$-order interactions ($p\leq q$). (b) Conceptual illustrations of our divide-and-conquer procedure for dealing with the dependence of renormalization flows on the ergodicity. (c) A system with $500$ units is defined, where pairwise interactions follow a Barabási-Albert network ($c=2$). We use this system to illustrate the selection of time scale $\tau^{\left(q\right)}$ based on the specific heat ($q=2$).
@@
    We find the eigenvalues of $\mathbf{L}^{\left(q\right)}_{k}$ that are smaller than $\frac{1}{\widehat{\tau}^{\left(q\right)}}$, where $\widehat{\tau}^{\left(q\right)}$ is the time scale for $q$-order simplicial complexes (see Sec. IV.2 for the selection of $\widehat{\tau}^{\left(q\right)}$). These eigenvalues correspond to long-range high-order interactions because the contributions of their eigenvectors have slow decays in Eq. (10). We denote the number of these eigenvalues as $m_{k}^{\left(q\right)}$. Other eigenvalues correspond to short-range high-order interactions and can be coarse grained. Moreover, according to Eq. (11), the totality of long-range high-order interactions (i.e., corresponding to smaller eigenvalues) is generally greater than that of short-range ones (i.e., associated with larger eigenvalues). Therefore, integrating out short-range high-order interactions does not damage the effective information contained in the system significantly.
@@
    In the moment space, we reduce the modes with fast decays in the system. The current $q$-order Laplacian, $\mathbf{L}^{\left(q\right)}_{k}$, is reduced to the re-scaled contributions of the eigenvectors associated with long-range high-order interactions
    $\mathbf{Q}^{\left(q\right)}_{k}=\sum_{\lambda^{\left(q\right)}_{k}}\Theta\left(\frac{1}{\widehat{\tau}^{\left(q\right)}}-\lambda^{\left(q\right)}_{k}\right)\lambda^{\left(q\right)}_{k}\big{|}\lambda^{\left(q\right)}_{k}\big{\rangle}\big{\langle}\lambda^{\left(q\right)}_{k}\big{|},$ (17)
@@
    where $\Theta\left(\cdot\right)$ is the Heaviside step function (i.e., $\Theta\left(x\right)=1$ if $x\geq 0$ and $\Theta\left(x\right)=0$ if $x<0$). Meanwhile, the current $p$-order Laplacian, $\mathbf{L}^{\left(p\right)}_{k}$, is reduced to the contributions of the $m_{k}^{\left(q\right)}$ smallest eigenvalues (i.e., long-range $p$-order interactions)
@@
    where $\Omega^{\left(p\right)}_{k}$ is the set of the $m_{k}^{\left(q\right)}$ smallest eigenvalues of $\mathbf{L}^{\left(p\right)}_{k}$ and $I_{A}\left(\cdot\right)$ is the indicator function defined on an arbitrary set $A$. Note that the eigenvalues in $\Omega^{\left(p\right)}_{k}$ are not necessarily smaller than $\frac{1}{\tau^{\left(q\right)}}$.
@@
    In the real space, we search for every cluster of units sharing short-range high-order interactions within the time scale $\widehat{\tau}^{\left(q\right)}$, such that we can coarse grain these units into a macro-unit. To this end, we generate a reference network, $\mathbf{H}^{\left(q\right)}_{k}$, in the following way. We first initialize $\mathbf{H}^{\left(q\right)}_{k}$ as a null network (i.e., with no edge). We follow Eq. (11) to calculate the density operator associated with time scale $\widehat{\tau}^{\left(q\right)}$ as $\widehat{\rho}^{\left(q\right)}_{k}=\exp\left(-\widehat{\tau}^{\left(q\right)}\mathbf{L}^{\left(q\right)}_{k}\right)/\operatorname{tr}\left(\exp\left(-\widehat{\tau}^{\left(q\right)}\mathbf{L}^{\left(q\right)}_{k}\right)\right)$. In operator $\widehat{\rho}^{\left(q\right)}_{k}$, its main diagonal elements, related to self-interactions of units, are frequently greater than most of its off-diagonal elements (e.g., see Fig. 2(b)). If the $\left(i,j\right)$-th element is greater than the $\left(i,i\right)$-th and $\left(j,j\right)$-th elements, then the $q$-order interactions between units $i$ and $j$ are sufficiently quick such that their accumulations within time scale $\widehat{\tau}^{\left(q\right)}$ are faster than those of self-interactions. In this case, we add an edge between units $i$ and $j$ in network $\mathbf{H}^{\left(q\right)}_{k}$. By progressively implement the edge adding procedure for every pair of units, we can form $n_{k}^{\left(q\right)}$ clusters in $\mathbf{H}^{\left(q\right)}_{k}$. Each cluster contains a set of strongly correlated units with short-range high-order interactions. In other words, these clusters can be understood as the Kadanoff blocks (i.e., the quasi-independent blocks whose mean size serves as the correlation length of the system) in the real space renormalization procedure [71, 72, 73]. In $\mathbf{G}^{\left(q\right)}_{k}$, each set of units that belong to the same cluster in $\mathbf{H}^{\left(q\right)}_{k}$ are aggregated into a super-unit to generate a new network $\mathbf{G}^{\left(q\right)}_{k+1}$. Consequently, there remain $n^{\left(q\right)}_{k}$ super-units in $\mathbf{G}^{\left(q\right)}_{k+1}$ after coarse graining. Parallel to this process, we also aggregate the units belonging to the same cluster of $\mathbf{H}^{\left(q\right)}_{k}$ into a super-unit in $\mathbf{G}^{\left(p\right)}_{k}$. The generated $\mathbf{G}^{\left(p\right)}_{k+1}$ contains $n^{\left(q\right)}_{k}$ super-units as well.
  • (5)
    The correspondence between the moment space and the real space for $q$-order interactions is ensured by a Laplacian $\mathbf{L}^{\left(q\right)}_{k+1}$. This operator is generated following $\left(\mathbf{L}^{\left(q\right)}_{k+1}\right)_{\nu_{i}\nu_{j}}=\big{\langle}\nu_{i}\big{|}\mathbf{Q}^{\left(q\right)}_{k}\big{|}\nu_{j}\big{\rangle}$, where $\nu_{i}$ and $\nu_{j}$ are super-units in $\mathbf{G}^{\left(q\right)}_{k+1}$. Notion $\big{|}\nu_{j}\big{\rangle}$ is a $N^{\left(q\right)}_{k}$-dimensional ket where unitary components correspond to all the units in $\mathbf{G}^{\left(q\right)}_{k}$ that are aggregated in $\nu_{i}$ and zero components correspond to all other units in $\mathbf{G}^{\left(q\right)}_{k+1}$ that are not covered by $\nu_{i}$. Mathematically, this procedure actually defines a similarity transformation $T$ between $\mathbf{Q}^{\left(q\right)}_{k}$ and $\mathbf{L}^{\left(q\right)}_{k+1}$
@@
    where $\mathbf{0}$ is a $\left(N^{\left(q\right)}_{k}-n^{\left(q\right)}_{k}\right)$-dimensional all-zero square matrix. It is clear that the similarity transformation is given as $T=\left(\big{|}\nu_{1}\big{\rangle},\ldots,\big{|}\nu_{n^{\left(q\right)}_{k}}\big{\rangle},\big{|}\eta_{n^{\left(q\right)}_{k}+1}\big{\rangle},\ldots,\big{|}\eta_{N^{\left(q\right)}_{k}}\big{\rangle}\right)$, where each ket $\eta_{j}$ is selected such that the columns of $T$ define a group of orthonormal bases. Similarly, we can derive the Laplacian $\mathbf{L}^{\left(p\right)}_{k+1}$ for $p$-order interactions based on $\mathbf{Q}^{\left(p\right)}_{k}$. The derived $\mathbf{L}^{\left(p\right)}_{k+1}$ and $\mathbf{L}^{\left(q\right)}_{k+1}$ are used in the $\left(k+1\right)$-th iteration.
The renormalization of $p$-order interactions in both the real and moment spaces is realized under the guidance of $q$-order interactions (e.g., see steps (3-4) for details). After renormalizing $\mathbf{L}^{\left(p\right)}_{k}$ by repeating steps (1-5), the SRG progressively drives the system towards an intrinsic scale of $p$-order interactions that exceeds the microscopic scales. More specifically, the SRG creates a $q$-order-guided iterative scale transformation that, irrespective of starting from which local scale, always leads to the latent critical point associated with $p$-order interactions (if there is any). It provides opportunities to verify whether the concerned thermodynamic functions are scale-invariant or not. See Fig. 3(a) for a summary of key steps in our approach. Note that the using of a reference network in step (4) is not necessary for theoretical derivations, yet it is favorable for computer programming.
@@
Figure 4: Instances of the renormalization flows of the simplex renormalization group (SRG). The first three iterations of renormalization on the systems whose pairwise interactions follow the random tree (a), the square lattice (b), the triangular lattice (c), and the Watts-Strogatz networks (d-e) are presented, where we set $\left(p,q\right)=\left(1,1\right)$, $\left(p,q\right)=\left(1,2\right)$ and $\left(p,q\right)=\left(1,3\right)$ for (a-b), (c-d) and (e), respectively.
@@
    If $\mathbf{G}^{\left(q\right)}_{k}$ is connected (i.e., the ergodicity holds), we directly apply steps (1-5) on $\mathbf{G}^{\left(q\right)}_{k}$ to guide the renormalization of $p$-order interactions and derive the inputs of the $\left(k+1\right)$-iteration.
@@
    If $\mathbf{G}^{\left(q\right)}_{k}$ is disconnected (i.e., the ergodicity does not hold) and has $r$ clusters denoted by $C_{1},\;\ldots,C_{r}$, we respectively apply steps (1-5) on each cluster of $\mathbf{G}^{\left(q\right)}_{k}$ to guide the renormalization of $p$-order interactions formed among the units that belong to this cluster. Specifically, we treat each cluster $C_{i}$ as a network and input it into steps (1-5) with $\mathbf{G}^{\left(p\right)}_{k}$. Please note that the above procedure does not affect $\mathbf{G}^{\left(p\right)}_{k}$ if the selected cluster $C_{i}$ is trivial (i.e., contains only one unit). After dealing with all $r$ clusters, the obtained results are used for the $\left(k+1\right)$-iteration.
@@
In Appendix B, we explain how the SRG is related to the persistent homology theory [62, 74, 75, 76] and suggest its potential applicability in analyzing the lifetime of topological properties. In Appendix C, we discuss the SRG in a form of conventional renormalization groups, suggesting the benefits of including high-order interactions into renormalization group analysis and indicate the relations between the SRG and other frameworks [31, 77, 78, 79]. Specifically, we find that the ideas underlying the SRG can be clearly interpreted by relating the SRG with a PCA-like renormalization group theory [31].

### v7

@@ III Simplex path integral
where $\operatorname{tr}\left(\cdot\right)$ denotes the trace and $\tau$ serves as the inverse of a finite temperature. Please see Fig. 2(b) for instances. The Boltzmann constant is assumed as unitary for simplicity. The partition function, $\sum_{\lambda^{\left(n\right)}}\exp\left(-\tau\lambda^{\left(n\right)}\right)$, is proportional to the average return probability of random walks within a time scale $\tau$ [42]. See Figs. 2(b-c) for illustrations.
@@
where $\hbar$ denotes the reduced Planck constant [3, 4]. Eq. (14) is closely related to the imaginary time evolution governed by operator $\mathbf{L}^{\left(n\right)}$ in the infinite-dimensional path space. Specifically, the probability amplitude for the system to evolve from $\big{|}\mathbf{x}_{0}\big{\rangle}$ to $\big{|}\mathbf{x}\big{\rangle}$ during a period of $\tau$ conveyed by the propagator in Eq. (13) can be measured by summing over all possible evolution paths connecting between these two states (i.e., the discrete counterpart of path integrals) in Eq. (14). This is the reason why the $n$-order network topology conveyed by $\mathbf{L}^{\left(n\right)}$ intrinsically shapes path integrals. See Fig. 2(c) for illustrations.
@@ IV Simplex renormalization group
The connection between Eqs. (8-16) and the moment space naturally inspires us to consider the generalization of renormalization group theories. The freedom degrees corresponding to short-range interactions (e.g., fast diffusion within the clusters of strongly correlated units) can be safely coarse grained without significant information loss, which is consistent with our interpretation of Eq. (10). Here we develop a framework to generalize renormalization group to high-order interactions, where we also propose possible ways to overcome the limitations of the Laplacian renormalization group (LRG) [10].
@@ IV.1 Renormalization procedure
Figure 3: Illustrations of the simplex renormalization group (SRG). (a) The conceptual illustration of the SRG is presented. For the sake of legibility, the simplicial complexes of different orders are not elaborated and are abstractly represented by dashed lines with different colors. Given a system where multi-order interactions coexist, the SRG can be applied to renormalize the system on the $p$-order based on the structure and dynamics associated with $q$-order interactions ($p\leq q$). (b) Conceptual illustrations of our divide-and-conquer procedure for dealing with the dependence of renormalization flows on the ergodicity. (c) A system with $500$ units is defined, where pairwise interactions follow a Barabási-Albert network ($c=2$). We use this system to illustrate the selection of time scale based on the specific heat ($q=2$).
@@
    In the real space, we first search targets to coarse grain. Specifically, we search for every cluster of units sharing short-range high-order interactions within the time scale $\widehat{\tau}^{\left(q\right)}$, such that we can coarse grain these units into a macro-unit. To this end, we generate a reference network, $\mathbf{H}^{\left(q\right)}_{k}$, in the following way. We first initialize $\mathbf{H}^{\left(q\right)}_{k}$ as a null network (i.e., with no edge). We follow Eq. (11) to calculate the density operator associated with time scale $\widehat{\tau}^{\left(q\right)}$ as $\widehat{\rho}^{\left(q\right)}_{k}=\exp\left(-\widehat{\tau}^{\left(q\right)}\mathbf{L}^{\left(q\right)}_{k}\right)/\operatorname{tr}\left(\exp\left(-\widehat{\tau}^{\left(q\right)}\mathbf{L}^{\left(q\right)}_{k}\right)\right)$. In operator $\widehat{\rho}^{\left(q\right)}_{k}$, its main diagonal elements, related to self-interactions of units, are frequently greater than most of its off-diagonal elements (e.g., see Fig. 2(b)). If the $\left(i,j\right)$-th element is greater than the $\left(i,i\right)$-th or the $\left(j,j\right)$-th element, then the $q$-order interactions between units $v_{i}$ and $v_{j}$ are sufficiently quick such that their accumulations within time scale $\widehat{\tau}^{\left(q\right)}$ are faster than those of self-interactions. In this case, we add an edge between units $v_{i}$ and $v_{j}$ in network $\mathbf{H}^{\left(q\right)}_{k}$. By progressively implement the edge adding procedure for every pair of units, we can form $n_{k}^{\left(q\right)}$ clusters in $\mathbf{H}^{\left(q\right)}_{k}$. Each cluster contains a set of strongly correlated units with short-range high-order interactions. In other words, these clusters can be understood as the Kadanoff blocks (i.e., the quasi-independent blocks whose mean size serves as the correlation length of the system) in the real space renormalization procedure [71, 72, 73].
@@
    After we find the targets for coarse graining in the real space, we implement a reormalization procedure. Specifically, in $\mathbf{G}^{\left(q\right)}_{k}$, each set of units that belong to the same cluster in $\mathbf{H}^{\left(q\right)}_{k}$ are aggregated into a macro-unit to generate a new network $\mathbf{G}^{\left(q\right)}_{k+1}$. Consequently, there remain $n^{\left(q\right)}_{k}$ macro-units in $\mathbf{G}^{\left(q\right)}_{k+1}$ after coarse graining. In $\mathbf{G}^{\left(q\right)}_{k+1}$, two macro-units, $v_{i}$ and $v_{j}$, are connected if at least one unit aggregated into $v_{i}$ is connected with one unit aggregated into $v_{j}$ in $\mathbf{G}^{\left(q\right)}_{k}$. If a unit or an interaction in $\mathbf{G}^{\left(q\right)}_{k}$ does not participate in any coarse graining according to $\mathbf{H}^{\left(q\right)}_{k}$, it is adopted into $\mathbf{G}^{\left(q\right)}_{k+1}$ without modification. Parallel to this process, we also follow the same rules to aggregate the units in $\mathbf{G}^{\left(p\right)}_{k}$ that belong to the same cluster of $\mathbf{H}^{\left(q\right)}_{k}$ into a macro-unit in $\mathbf{G}^{\left(p\right)}_{k+1}$. The generated $\mathbf{G}^{\left(p\right)}_{k+1}$ contains $n^{\left(q\right)}_{k}$ macro-units as well. Therefore, we define $N^{\left(p\right)}_{k+1}=N^{\left(q\right)}_{k+1}=n^{\left(q\right)}_{k}$.
  • (4)
    In the moment space, we first search modes to reduce. Specifically, we find the eigenvalues of $\mathbf{L}^{\left(q\right)}_{k}$ that are smaller than $\frac{1}{\widehat{\tau}^{\left(q\right)}}$, where $\widehat{\tau}^{\left(q\right)}$ is the time scale for $q$-order simplicial complexes (see Sec. IV.2 for the selection of $\widehat{\tau}^{\left(q\right)}$). These eigenvalues correspond to long-range high-order interactions because the contributions of their eigenvectors have slow decays in Eq. (10). We denote the number of these eigenvalues as $s_{k}^{\left(q\right)}$. Other eigenvalues correspond to short-range high-order interactions and can be coarse grained. Moreover, according to Eq. (11), the totality of long-range high-order interactions (i.e., corresponding to smaller eigenvalues) is generally greater than that of short-range ones (i.e., associated with larger eigenvalues). Therefore, integrating out short-range high-order interactions does not damage the effective information contained in the system significantly. The only problem is that the eigenvalue spectrum cutoff defined above may overestimate the number of short-range high-order interactions and make $s_{k}^{\left(q\right)}$ too small. This is because the eigenvalue spectrum of $\mathbf{L}^{\left(q\right)}_{k}$ does not reflect a possibility that the short-range $q$-order interactions between units $v_{i}$ and $v_{j}$ can be slower than their self-interactions. For instance, although the interactions between units $v_{i}$ and $v_{j}$ are classified as short-range, the $\left(i,j\right)$-th element of $\widehat{\rho}^{\left(q\right)}_{k}$ can be smaller than both the $\left(i,i\right)$-th and the $\left(j,j\right)$-th elements. In this case, the time scale $\widehat{\tau}^{\left(q\right)}$ may be too short for the correlations between $v_{i}$ and $v_{j}$ to accumulate to a non-negligible level. Compared with self-interactions, the interactions between units $v_{i}$ and $v_{j}$ are less important in determining the behaviours of $v_{i}$ and $v_{j}$ since these interactions are not strong enough. To avoid the possible overestimation, we use $n_{k}^{\left(q\right)}$, the number of Kadanoff blocks in $\mathbf{H}^{\left(q\right)}_{k}$, as a bound. In the case with overestimation, the number of long-range interactions is underestimated (i.e., $s_{k}^{\left(q\right)}\leq n_{k}^{\left(q\right)}$) and we correct it by defining $m_{k}^{\left(q\right)}=\max\{s_{k}^{\left(q\right)},n_{k}^{\left(q\right)}\}$. The value of $m_{k}^{\left(q\right)}$ defines the number of modes to keep during renormalization.
  • (5)
@@
    After we find the modes to reduce in the moment space, we exclude these modes with fast decays from the system such that there remain $m_{k}^{\left(q\right)}$ modes. Specifically, the current $q$-order Laplacian, $\mathbf{L}^{\left(q\right)}_{k}$, is reduced to the re-scaled contributions of the eigenvectors associated with long-range high-order interactions
    $\mathbf{Q}^{\left(q\right)}_{k}=\sum_{\lambda^{\left(q\right)}_{k}}I_{\Lambda^{\left(q\right)}_{k}}\left(\lambda^{\left(q\right)}_{k}\right)\lambda^{\left(q\right)}_{k}\big{|}\lambda^{\left(q\right)}_{k}\big{\rangle}\big{\langle}\lambda^{\left(q\right)}_{k}\big{|},$ (17)
@@
    where $\Lambda^{\left(q\right)}_{k}$ is the set of the $m_{k}^{\left(q\right)}$ smallest eigenvalues of $\mathbf{L}^{\left(q\right)}_{k}$ (i.e., the first $m_{k}^{\left(q\right)}$ eigenvalues after we sort the eigenvalue spectrum in a decreasing order) and $I_{A}\left(\cdot\right)$ is the indicator function defined on an arbitrary set $A$. Meanwhile, the current $p$-order Laplacian, $\mathbf{L}^{\left(p\right)}_{k}$, is reduced to the contributions of the $m_{k}^{\left(q\right)}$ smallest eigenvalues (i.e., long-range $p$-order interactions)
@@
    where $\Omega^{\left(p\right)}_{k}$ is the set of the $m_{k}^{\left(q\right)}$ smallest eigenvalues of $\mathbf{L}^{\left(p\right)}_{k}$. Note that the eigenvalues in $\Omega^{\left(p\right)}_{k}$ are not necessarily smaller than $\frac{1}{\widehat{\tau}^{\left(q\right)}}$.
@@
    The correspondence between moment space and real space for $q$-order interactions is ensured by a Laplacian $\mathbf{L}^{\left(q\right)}_{k+1}$. For $v_{i}$ and $v_{j}$, two macro-units in $\mathbf{G}^{\left(q\right)}_{k+1}$, we define
    $\left[\mathbf{L}^{\left(q\right)}_{k+1}\right]_{v_{i}v_{j}}=\big{\langle}v_{i}\big{|}\mathbf{Q}^{\left(q\right)}_{k}\big{|}v_{j}\big{\rangle},$ (19)
    if $v_{i}$ and $v_{j}$ are connected in $\mathbf{G}^{\left(q\right)}_{k+1}$. We set $\left[\mathbf{L}^{\left(q\right)}_{k+1}\right]_{v_{i}v_{j}}=0$ if $v_{i}$ and $v_{j}$ are disconnected in $\mathbf{G}^{\left(q\right)}_{k+1}$. Here $\big{|}v_{j}\big{\rangle}$ is a $N^{\left(q\right)}_{k}$-dimensional ket where unitary components correspond to all the units in $\mathbf{G}^{\left(q\right)}_{k}$ that are aggregated into $v_{i}$ and zero components correspond to all other units in $\mathbf{G}^{\left(q\right)}_{k+1}$ that are not covered by $\nu_{i}$ (i.e., the value of $\left[\mathbf{L}^{\left(q\right)}_{k+1}\right]_{v_{i}v_{j}}$ is the sum of all $\left[\mathbf{Q}^{\left(q\right)}_{k}\right]_{xy}$ if $v_{i}$ and $v_{j}$ are connected, where $x$ goes through all the units in $\mathbf{G}^{\left(q\right)}_{k}$ that are aggregated into $v_{i}$ and $y$ goes through all units in $\mathbf{G}^{\left(q\right)}_{k}$ that are aggregated into $v_{j}$). After defining all off-diagonal elements, the diagonal elements of $\mathbf{L}^{\left(q\right)}_{k+1}$ are defined following
    $\left[\mathbf{L}^{\left(q\right)}_{k+1}\right]_{v_{i}v_{i}}=-\sum_{v_{j}}\left[\mathbf{L}^{\left(q\right)}_{k+1}\right]_{v_{i}v_{j}}.$ (20)
    Mathematically, this procedure actually defines a similarity transformation $T$ between $\mathbf{Q}^{\left(q\right)}_{k}$ and $\mathbf{L}^{\left(q\right)}_{k+1}$
@@
    where $\mathbf{0}$ is a $\left(N^{\left(q\right)}_{k}-n^{\left(q\right)}_{k}\right)$-dimensional all-zero square matrix. It is clear that the similarity transformation is given as $T=\left(\big{|}v_{1}\big{\rangle},\ldots,\big{|}v_{n^{\left(q\right)}_{k}}\big{\rangle},\big{|}\eta_{n^{\left(q\right)}_{k}+1}\big{\rangle},\ldots,\big{|}\eta_{N^{\left(q\right)}_{k}}\big{\rangle}\right)$, where each ket $\eta_{j}$ is selected such that the columns of $T$ define a group of orthonormal bases. Similarly, we can derive the Laplacian $\mathbf{L}^{\left(p\right)}_{k+1}$ for $p$-order interactions based on $\mathbf{Q}^{\left(p\right)}_{k}$. The derived $\mathbf{L}^{\left(p\right)}_{k+1}$ and $\mathbf{L}^{\left(q\right)}_{k+1}$ are used in the $\left(k+1\right)$-th iteration.
The renormalization of $p$-order interactions in both the real space (i.e., see steps (2-3)) and the moment space (i.e., see steps (4-5)) is realized under the guidance of $q$-order interactions. After renormalizing $\mathbf{L}^{\left(p\right)}_{k}$ by repeating steps (1-6), the SRG progressively drives the system towards an intrinsic scale of $p$-order interactions that exceeds the microscopic scales. More specifically, the SRG creates a $q$-order-guided iterative scale transformation that, irrespective of starting from which local scale, always leads to the latent critical point associated with $p$-order interactions (if there is any). It provides opportunities to verify whether the concerned thermodynamic functions are scale-invariant or not. See Fig. 3(a) for a summary of key steps in our approach. Note that the using of a reference network in step (2) is not necessary for theoretical derivations, yet it is favorable for computer programming.
@@
Figure 4: Illustrations of the iteration process and coarse graining procedure in the simplex renormalization group (SRG). (a) summarizes key processing pipelines of the SRG proposed in Secs. IV.1-IV.2. Specifically, given an input system and a selected combination of $\left(p,q\right)$, we follow step (1) to characterize the system on the $p$- and $q$-orders to generate its representations in the real (i.e., $\mathbf{G}^{\left(p\right)}_{1}$ and $\mathbf{G}^{\left(q\right)}_{1}$) and the moment (i.e., $\mathbf{L}^{\left(p\right)}_{1}$ and $\mathbf{L}^{\left(q\right)}_{1}$) spaces. Then, we follow Sec. IV.2 to calculate the time scale $\widehat{\tau}^{\left(q\right)}$. If the ergodicity does not holds in $\mathbf{G}^{\left(q\right)}_{1}$, we repeat the time scale calculation on each cluster of $\mathbf{G}^{\left(q\right)}_{1}$. After obtaining $\widehat{\tau}^{\left(q\right)}$, we follow steps (2-3) to find the Kadanoff blocks and implement coarse graining in the real space for generating $\mathbf{G}^{\left(p\right)}_{2}$ and $\mathbf{G}^{\left(q\right)}_{2}$. We also follow steps (4-5) to find the modes corresponding to short-range high-order interactions and reduce the system in the moment space into the contributions of long-range high-order interactions. Finally, we calculate $\mathbf{L}^{\left(p\right)}_{2}$ and $\mathbf{L}^{\left(q\right)}_{2}$ according to the correspondence between moment space and real space. The generated $\mathbf{G}^{\left(p\right)}_{2}$, $\mathbf{G}^{\left(q\right)}_{2}$, $\mathbf{L}^{\left(p\right)}_{2}$, and $\mathbf{L}^{\left(q\right)}_{2}$ are used for the next iteration of the SRG. (b-d) use a system whose pairwise interactions follow the Watts-Strogatz network to show the real space renormalization flows on $1$-order when the SRG is guided by $q\in\{1,2,3\}$, respectively. In these instances, the SRG runs two iterations and we assign each macro-unit in $G^{\left(q\right)}_{3}$ with a unique color. If a unit is aggregated into a certain macro-unit, it adopts the color from this macro-unit. By recursively repeat this procedure, we can assign units in each iteration with appropriate colors to indicate which macro-unit in $G^{\left(q\right)}_{3}$ contains them. Meanwhile, to highlight the effects of different values of $q$ on renormalization flows, we assign each unit in $G^{\left(q\right)}_{1}$ with an index to trace the units grouped into the Kadanoff blocks in the step (2) of the SRG framework. For instance (see (c)), the units $1$, $2$, $3$, $28$, $29$, and $30$ in $G^{\left(1\right)}_{1}$ are aggregated into a macro-unit when $q=2$, whose index is assigned as $1$ in $G^{\left(1\right)}_{2}$. Meanwhile, the units $10$, $11$, $12$, $13$, and $14$ in $G^{\left(1\right)}_{1}$ are coarse grained into a macro-unit in $G^{\left(1\right)}_{2}$, which is indexed as $10$. Moreover, the units $17$, $18$, $20$, and $15$ in $G^{\left(1\right)}_{1}$ are grouped into the macro-unit $17$ in $G^{\left(1\right)}_{2}$. Therefore, the system size is reduced from $30$ to $18$ after this time of renormalization.
@@
    If $\mathbf{G}^{\left(q\right)}_{k}$ is connected (i.e., the ergodicity holds), we directly apply steps (1-6) on $\mathbf{G}^{\left(q\right)}_{k}$ to guide the renormalization of $p$-order interactions and derive the inputs of the $\left(k+1\right)$-iteration.
@@
    If $\mathbf{G}^{\left(q\right)}_{k}$ is disconnected (i.e., the ergodicity does not hold) and has $r$ clusters denoted by $C_{1},\;\ldots,C_{r}$, we respectively apply steps (1-6) on each cluster of $\mathbf{G}^{\left(q\right)}_{k}$ to guide the renormalization of $p$-order interactions formed among the units that belong to this cluster. Specifically, we treat each cluster $C_{i}$ as a network and input it into steps (1-6) with $\mathbf{G}^{\left(p\right)}_{k}$. Please note that the above procedure does not affect $\mathbf{G}^{\left(p\right)}_{k}$ if the selected cluster $C_{i}$ is trivial (i.e., contains only one unit). After dealing with all $r$ clusters, the obtained results are used for the $\left(k+1\right)$-iteration.
@@
In Appendix C, we explain how the SRG is related to the persistent homology theory [62, 74, 75, 76] and suggest its potential applicability in analyzing the lifetime of topological properties. In Appendix D, we discuss the SRG in a form of conventional renormalization group, suggesting the benefits of including high-order interactions into renormalization group analysis and indicate the relations between the SRG and other frameworks [31, 77, 78, 79]. Specifically, we find that the ideas underlying the SRG can be clearly interpreted by relating the SRG with a PCA-like renormalization group theory [31].